D3-3D²D'+ D'³ z= 0
1
Expert's answer
2020-07-08T19:41:33-0400

Given "(D^3-3D\u00b2D'+ D'\u00b3) z= 0" where "D = \frac{\partial}{\partial x}, D' = \frac{\partial}{\partial y}" .

Since, "D^3-3D\u00b2D'+ D'\u00b3" is irreducible in linear factors,

assume solution is "z = \sum A e^{hx+ky}" .

Now. "(D^3-3D²D'+ D'³)z =(h^3-3h^2k+k^3) \sum{Ae^{hx+ky}} = 0" "\implies h^3 - 3h^2 k +k^3 = 0" .

Hence, "z = \sum A e^{hx+ky}" is solution of given differential equation where h and k holds the condition "h^3 - 3h^2 k +k^3 =0" .